OpenStudy (anonymous):

using left hand endpoints with 2 subintervals, approximate integral [0,1] (1/1+x^2)dx

OpenStudy (anonymous):

\[\int\limits_{0}^{1} 1\div \left( 1+x^{2} \right) dx\]

OpenStudy (bahrom7893):

oh man i really don't feel like doin any math right now.. but seems like u need help and nobody's helpin.. i guess ill try this..

OpenStudy (anonymous):

do you have the selection answers?

OpenStudy (bahrom7893):

Okay so first of all it says two sub intervals right ..and your integral is from 0 to 1, or the length is 1. So each of the steps will be: 1 / 2 = 0.5

OpenStudy (bahrom7893):

\[x _{0}=0\] since we are starting from 0 so: \[x _{0}=0\] \[x _{1}=0.5\] \[x _{2}=1\]

OpenStudy (bahrom7893):

As you can see, each step is the previous one plus 0.5

OpenStudy (bahrom7893):

1) \[f(x _{0}) = 1/(1+0^2) = 1\]

OpenStudy (bahrom7893):

2) \[f(x _{1}) = 1/(1+0.5^2)\]

OpenStudy (bahrom7893):

=0.8

OpenStudy (bahrom7893):

3) \[f(x_2)=1/(1+1^2) = 1/2\]

OpenStudy (bahrom7893):

Now to approximate the integral: use: \[h*(f(x_0)+f(x_1)+f(x_2))\]

OpenStudy (bahrom7893):

where h is the step size or 0.5

OpenStudy (bahrom7893):

So you end up with: 0.5 ( 1 + 0.8 + 0.5 ) = 0.5 * 2.3 = 1.15

OpenStudy (bahrom7893):

The actual answer is 0.785 or pi/4, but we are so off because we only have 2 steps.. the more steps you have or the less the step size, the better is ur answer.

OpenStudy (anonymous):

ah we are looking for the LH... n=2, deltax=(1-0)/2=1/2=.5 LH=(1/2)(1+.8)=.9 answer and for the RH=(1/2)(.8+.5)=13/20=.65

OpenStudy (bahrom7893):

mark you are wrong.. you didn't add the last step.. you did from 0 to 1/2

OpenStudy (bahrom7893):

then he needs 1/2 to 1.. which will be: (1/2) ( 1 + .8 + .5) = 1.15

OpenStudy (anonymous):

nope,, the LH rule is up to .5only, and for the RH rule is up to 1

OpenStudy (anonymous):

n= 0 ,5 1 f(x)= 1 .8 .5

OpenStudy (bahrom7893):

http://www.mathsnetalevel.com/question.php?ref=3805

OpenStudy (bahrom7893):

follow the steps and see how it goes..

OpenStudy (anonymous):

i think i got it from here just needed a refresher. thanks guys!

OpenStudy (anonymous):

LH=(delta x)(f(0)+f(.5))=.5(1+.8)= .9 and for the RH RH=(deltax)(f(.5)+f(1))=.5(.8+.5)=.65

OpenStudy (anonymous):

ok have fun guys......thnx